# It’s Morphing Time: Unleashing the Potential of Multiple LLMs via Multi-objective Optimization

### v2

@@ Abstract
In this paper, we introduce a novel approach for large language model merging via black-box multi-objective optimization algorithms. The goal of model merging is to combine multiple models, each excelling in different tasks, into a single model that outperforms any of the individual source models. However, model merging faces two significant challenges: First, existing methods rely heavily on human intuition and customized strategies to tackle multiple tasks. Second, it’s difficult to search for the great model merging configuration in limited evaluations. To address these challenges, we propose a multi-objective optimization based model merging method named MM-MO. The proposed method can automatically search merging configurations for multiple tasks with multi-objective optimization algorithms. Moreover, to obtain high-quality model merging configurations within a limited number of evaluation iterations, we have made several improvements to multi-objective Bayesian optimization specifically for model merging scenarios. First, we introduced a weak-to-strong method to improve the acquisition strategy. Second, we employed Fisher information to select configurations, further increasing the chances of discovering superior model merging configurations. Third, we designed a sparsity metric as an additional optimization objective to enhance the model’s generalization performance across different tasks. We conducted comprehensive experiments with other mainstream model merging methods, demonstrating that our method consistently outperforms them. Moreover, performance improvements are observed even on the tasks not explicitly targeted as optimization objectives, indicating that our method enhances the overall potential of the model. The proposed method provides a significant advancement in model merging techniques, offering a robust and efficient solution for integrating diverse models into a unified, high-performing model without additional training data.
1 Introduction
@@
Large language models (LLMs) have shown great performance on tasks in various domains such as natural language processing [1], computer vision [2] etc. With the great effort of world-wide contributors in community, a large number of general-purpose pre-trained and task-specific fine-tuned language models have been proposed and made publicly available. However, LLM pre-training or fine-tuning is non-trivial and requires a lot of effort and financial budget. Recently, model merging (MM) has attracted many researchers’ attention. By combining multiple LLMs into a single model with better performance and adaptability on more tasks, MM offers a novel cost-efficient way of obtaining new powerful language models without performing additional training [3, 4], just like the Power Rangers merge their Zords together to form a mighty Megazord [5]. Ideally, MM is supposed to inherit and amplify the strengths from its source models while ignoring their weaknesses. Therefore, the obtained model will be able to tackle the union set of all the tasks where the source models are pre-trained/fine-tuned with better performance. Yet this is achieved without training, which saves a large amount of calculation/financial budget. With the help of open-source toolkits such as mergekit [6, 7], MM has become popular for LLM developing and shown great potential on the Open LLM Leaderboard [8].
@@
However, model merging requires the model maker to have profound knowledge or intuition. Automatically discovering more capable MM recipes is still in its infancy. To the best of our knowledge, The most related work is [3], where Akiba et al. proposed to use evolutionary algorithms (EAs) to generate powerful model merging recipes which operates in both parameter space and data flow space. However, the approach of using diverse task scores and optimizing the parameter space to explore the comprehensive potential of the final model has been largely overlooked.
@@
MM is similar to ensemble learning in the sense that they both try to produce one model from multiple source models. In ensemble learning, It is generally believed that diversity in an ensemble could be beneficial for improving its performance [9, 10, 11]. One could infer that maintaining diversity during MM may also result in powerful LLMs. Having this in mind, we design a Model Merging approach via multi-objective optimization which take into consideration multiple task-specific performance metrics simultaneously.
@@
(1) Automated Model Merging via multi-objective optimization: We introduce a novel approach to model merging that automates the search for optimal merging configurations using multi-objective black-box optimization algorithms, eliminating the reliance on human intuition and customized strategies.
@@
(3) Additional Optimization Objective: In addition to task metrics, we have introduced a sparsity metric as an another optimization objective to enhance the model’s generalization performance across different tasks.
The remainder of this paper is organized as follows. Section 2 surveys the existing methods in model merging and multi-objective optimization. Section 3 introduces a novel multi-objective Bayesian optimization algorithm, MM-MO, tailored for model merging. In this section, we also present our improvements to the process of searching for new solutions and the design of new optimization objectives specifically for model merging scenarios. Section 4 presents the performance evaluation of the proposed method against related methods. Finally, we draw conclusions in Section 5.
2 Background
2.1 Model Merging
@@
Our approach builds upon the combination of TIES-Merging and DARE, leveraging their strengths to create a robust model merging strategy. By integrating the conflict resolution techniques of TIES-Merging with the parameter sparsification methods of DARE, we aim to automate the search for optimal model merging configurations, ensuring that the final merged model maintains high performance without the degradation typically associated with parameter conflicts.
2.2 Multi-objective Optimization
@@
Definition 2.1 (Pareto dominance [19])
@@
Multi-objective Optimization (MO) focuses on approximating the PS, targeting at a solution set with good convergence and diversity in the objective space. In this work, we utilize multi-objective optimization to address the complex challenge of model merging. By considering multiple objectives simultaneously, MO enables us to balance various performance metrics, such as different task-specific accuracies. This approach allows us to optimize the trade-offs between different performance criteria, leading to a more robust and effective merged model. Through this methodology, we aim to achieve a unified model that not only excels in specific tasks but also retains a broad and diverse set of capabilities from the individual models. This is particularly important in the context of LLMs, where different tasks may require different strengths and features from the base models. By leveraging MO, our method systematically explores the parameter space to identify optimal configurations, ensuring that the final model exhibits both high performance and comprehensive potential across various tasks.
3 Methodology
3.1 Overview
@@
In this section, we provide an overview of our proposed automated model merging approach, MM-MO. Our approach is designed to merge multiple pre-trained models into a single, unified model that retains the strengths of each individual model and surpasses the performance of any single model. Our automated approach aims at searching for the optimal merging configuration is discovered without relying on human intuition.
In Figure 1, we show the overall process of model merging based on multi-objective optimization, which is divided into three parts:
On the left is a schematic diagram of the automatic search process for model merging configurations. Our method starts with a set of source models, each obtained by performing SFT on different datasets using the same pre-trained backbone. These models have their own advantages for different task types. We merge these models to obtain a merged model, evaluate the performance of the merged model on different tasks as optimization objectives , and calculate the sparsity metric of the merged model as an additional optimization objective. For simplicity, we will denote $\boldsymbol{f}(\boldsymbol{x,Tj})$ later as $\boldsymbol{f_{j}}(\boldsymbol{x})$, $j\in\{1,...,M-1\}$. MM-MO generates new model merging configurations $\boldsymbol{X}=\{\boldsymbol{x}_{1},\boldsymbol{x}_{2},...\}$ based on these objective values, allowing the creation of new merged models. This process is repeated until a preset evaluation iteration limit is reached, ultimately resulting in an optimal model merging configuration. In the middle is the complete process of using multi-objective Bayesian optimization to search for model merging configurations, detailed in Section 3.2. On the right is the step of generating new configurations. We enhanced the acquisition strategy of multi-objective Bayesian optimization by designing a weak-to-strong method to adjust candidate configurations, and we used Fisher information to select the next generation of model merging configurations, detailed in Section 3.3. We also present the pseudocode of the MM-MO framework in the appendix (Section A).
3.2 Model Merging via Multi-objective Optimization
@@
Our approach builds upon TIES-Merging with DARE [17, 21], utilizing qEHVI to search for the optimal model merging configuration. The integration of qEHVI allows for an automated and systematic exploration of the parameter space, identifying the best configurations that balance various performance metrics. Here are the detailed steps of our methodology.
1) Initialization: We start with a collection of pre-trained models, each fine-tuned on different tasks. These models and their configurations serve as the initial candidates for merging.
2) Objective Definition: We define multiple objectives that reflect the performance metrics we aim to optimize. Specifically, these objectives include task-specific accuracies, such as those from C-EVAL [22] and GSM8K [23], as well as a sparsity metric $\boldsymbol{g}(\boldsymbol{x})$ used in our experiments. These metrics are crucial as they ensure the merged model performs well across different types of tasks.
3) Surrogate Model Training: qEHVI leverages surrogate models to approximate the objective functions. We train these surrogate models on the initial set of model configurations and their corresponding performance metrics. This step involves creating a probabilistic model that predicts the performance of unseen configurations.
4) Acquisition Function Optimization: The key component of qEHVI is the Expected Hypervolume Improvement (EHVI) acquisition function. This function guides the search by quantifying the expected improvement in the objective space. We optimize this acquisition function to select the most promising model configurations to evaluate next. Specifically, the hypervolume (HV) is a measure of the space dominated by a set of solutions in the objective space, bounded from below by a reference point. The HV indicator is defined as follows:
@@
where $\tilde{f}_{t}\sim p(\boldsymbol{f}|\mathcal{D})$ for $t=1,...,N$ and $\mathcal{X}_{\mathrm{cand}}=\{x_{i}\}_{i=1}^{q}$ [20]. After obtaining candidate configurations using the acquisition function, we adjusted the candidate configurations using the weak-to-strong method. We then selected the configurations based on their Fisher information to ultimately obtain a new generation of model merging configurations. For details, please refer to Section 3.3.
5) Parallel Evaluation: Unlike traditional methods, qEHVI supports parallel evaluations, which significantly speeds up the optimization process. In each iteration, we select a batch of model configurations and evaluate their performance on the defined objectives (e.g., C-EVAL and GSM8K accuracy).
6) Updating the Surrogate Model: After evaluating the selected configurations, we update the surrogate models with the new data. This iterative process continues, refining the surrogate models and improving the acquisition function’s accuracy in identifying high-potential configurations.
7) Convergence and Final Selection: The optimization process continues until convergence criteria are met, such as a maximum number of iterations or negligible improvement in the objective space. The final model merging configuration is selected based on the best trade-offs between the multiple performance metrics.
@@
3.3 Enhanced Acquisition Strategy
To achieve high-quality model merging configurations within a limited number of evaluation iterations, we have further improved the acquisition function of multi-objective Bayesian optimization for the model merging scenario.

### v3

@@ Abstract
In this paper, we introduce a novel approach for addressing the multi-objective optimization problem in large language model merging via black-box multi-objective optimization algorithms. The goal of model merging is to combine multiple models, each excelling in different tasks, into a single model that outperforms any of the individual source models. However, model merging faces two significant challenges: First, existing methods rely heavily on human knowledge or intuition. Second, it’s difficult to obtain the great model merging configuration in limited evaluations. To address these challenges, we formalize model merging as a multi-objective optimization problem and propose an automated optimization approach named MM-MO. This method leverages multi-objective optimization algorithms to autonomously search for optimal merging configurations across various tasks, alleviating the need for human intervention. In MM-MO, a weak-to-strong method is employed to enhance the acquisition function, allowing previously evaluated superior configurations to guide the search for new ones. Meanwhile, Fisher information is applied to screen these configurations, increasing the possibility of identifying high-quality merging configuration. Additionally, we designed a sparsity metric as an additional optimization objective to enhance the model’s generalization performance across different tasks. We conducted comprehensive experiments with other mainstream model merging methods, demonstrating that the proposed MM-MO algorithm is competitive and effective in achieving high-quality model merging.
Index Terms:
Large language model, model merging, multi-objective optimization
I Introduction
@@
Large language models (LLMs) have shown great performance on tasks in various domains such as natural language processing [1], computer vision [2] etc. With the great effort of world-wide contributors in community, a large number of general-purpose pre-trained and task-specific fine-tuned language models have been proposed and made publicly available [3, 4, 5, 6, 7, 8]. However, LLM pre-training or fine-tuning is non-trivial and requires a lot of effort and financial budget [9]. Recently, model merging (MM) has attracted many researchers’ attention. By combining multiple LLMs into a single model with better performance and adaptability on more tasks, MM offers a novel cost-efficient way of obtaining new powerful language models without requiring the collection of raw training data or expensive computation [10, 11, 12], just like the Power Rangers merge their Zords together to form a mighty Megazord [13]. Ideally, MM is supposed to inherit and amplify the strengths from its source models while ignoring their weaknesses. Therefore, the obtained model will be able to tackle the union set of all the tasks where the source models are pre-trained/fine-tuned with better performance. Yet this is achieved without training, which saves a large amount of calculation/financial budget. With the help of open-source toolkits such as mergekit [14, 15], MM has become popular for LLM developing and shown great potential on the Open LLM Leaderboard [16].
@@
However, model merging requires the model maker to have profound knowledge or intuition [11]. Alternatively, they may invest substantial time in conducting an extensive grid search to determine the configuration for model merging. For example, DARE [17] search the drop rate $p$ of each model in [0.1, 0.2, · · · , 0.9]. Since model merging involves not only determining the retention ratio of parameters but also setting the weight proportions of each model, the search space for a grid search expands substantially as the number of models increases. As a result, this approach becomes increasingly inefficient for identifying an optimal model merging configuration. Automatically and efficiently discovering more capable MM recipes is still in its infancy. To the best of our knowledge, The most related work is [11], where Akiba et al. proposed to use evolutionary algorithms (EAs) to generate powerful model merging recipes which operates in both parameter space and data flow space. However, the approach of using diverse task scores and optimizing the parameter space to explore the comprehensive potential of the final model has been largely overlooked.
@@
MM is similar to ensemble learning [18] in that both methods aim to leverage the strengths of multiple source models to improve overall performance across diverse tasks. In ensemble learning, It is generally believed that diversity in an ensemble could be beneficial for improving its performance [19, 20, 21]. One could infer that maintaining diversity during MM may also result in powerful LLMs. Having this in mind, we design a model merging approach via multi-objective optimization which take into consideration multiple task-specific performance metrics simultaneously.
Figure 1: Illustration of merged model performance differences when formalizing the model merging configuration search as a single-objective vs. multi-objective optimization problem. Multi-objective optimization achieves a more balanced performance across code and math capabilities.
In Fig 1, we illustrate the performance of merged models when the search for optimal model merging configurations is framed as either a single-objective or multi-objective optimization problem. When focusing solely on the model’s code or math capabilities, single-objective optimized configurations tend to concentrate at the extremes, typically enhancing one capability at the expense of the other. In contrast, multi-objective optimized configurations achieve a more balanced performance, resulting in a merged model that attains satisfactory levels in both Code and Math abilities. Given that the goal of model merging is to combine multiple models, each excelling in different tasks, into a single model that performs well across various tasks, approaching the configuration search as a multi-objective optimization problem is a more effective and appropriate solution.
@@
(1) Formalizing Model Merging as a Multi-Objective Optimization Problem: We introduce a novel perspective by formulating model merging as a multi-objective optimization problem. This approach allows for the simultaneous consideration of multiple task-specific performance metrics, aiming to creating a comprehensive model that integrates the strengths of multiple source models across diverse tasks.
(2) Automated Model Merging via multi-objective optimization: We introduce a novel approach to model merging that automates the search for better merging configurations using multi-objective black-box optimization algorithms, alleviating the reliance on human intuition and customized strategies.
@@
(4) Additional Optimization Objective: In addition to task metrics, we have introduced a sparsity metric as another optimization objective to enhance the model’s generalization performance across different tasks.
The remainder of this paper is organized as follows. Section II surveys the existing methods in model merging and multi-objective optimization. In Section III, we formalize model merging as a multi-objective optimization problem, establishing the framework that underpins our approach. Section IV introduces a novel multi-objective Bayesian optimization algorithm, MM-MO, tailored for model merging. In this section, we also present our improvements to the process of searching for new solutions and the design of new optimization objectives specifically for model merging scenarios. Section V and VI presents the performance evaluation of the proposed method against related methods. Finally, we draw conclusions in Section VII.
II Background
II-A Model Merging
@@
Our approach builds upon the combination of TIES-Merging and DARE, leveraging their strengths to create a robust model merging strategy. By integrating the conflict resolution techniques of TIES-Merging with the parameter sparsification methods of DARE, we aim to automate the search for better model merging configurations, ensuring that the final merged model maintains high performance without the degradation typically associated with parameter conflicts.
II-B Multi-objective Optimization
@@
Definition 1 (Pareto dominance [28])
@@
Multi-objective Optimization (MO) [29] focuses on approximating the PS, targeting at a solution set with good convergence and diversity in the objective space. In this work, we utilize multi-objective optimization to address the complex challenge of model merging. By considering multiple objectives simultaneously, MO enables us to balance various performance metrics, such as different task-specific accuracies. This approach allows us to optimize the trade-offs between different performance criteria, leading to a more robust and effective merged model. Through this methodology, we aim to achieve a unified model that not only excels in specific tasks but also retains a broad and diverse set of capabilities from the individual models. This is particularly important in the context of LLMs, where different tasks may require different strengths and features from the base models. By leveraging MO, our method systematically explores the parameter space to identify better configurations, ensuring that the final model exhibits both high performance and comprehensive potential across various tasks.
III Problem Formulation
Figure 2: An illustration of formalizing model merging as a multi-objective optimization problem, with three source models and three evaluation tasks.
Here, we formalize model merging as a multi-objective optimization problem. The goal is to combine multiple homologous source models into a single, unified model that retains the strengths of each individual model and achieves superior performance across diverse tasks compared to any individual source model. This formalization allows us to automate the search for an optimal merging configuration, eliminating the reliance on human intuition and custom strategies.
The problem is framed as follows: Given a set of source models ${M_{1},M_{2},…,M_{N}}$ (In Fig. 2, $N=3$), each of which is fine-tuned on distinct datasets but based on a common pre-trained backbone, our goal is to find a merging configuration that effectively balances task performance across multiple objectives. Each source model demonstrates strengths in different types of tasks, and the challenge is to derive a merged model that synthesizes these strengths for enhanced generalizability.
As illustrated in Fig. 2, the model merging process based on multi-objective optimization is divided into three main components:
1. Configuration Search for Model Merging: The process begins with the initialization of merging configurations $\boldsymbol{X}={\boldsymbol{x}_{1},\boldsymbol{x}_{2},…}$, where each configuration $\boldsymbol{x}_{i}$ represents a specific configuration for merging the source models. For each configuration, a merged model is generated and evaluated based on its performance on various tasks, represented as optimization objectives $\boldsymbol{f}(\boldsymbol{x,T1}),\boldsymbol{f}(\boldsymbol{x,T2}),\dots,\boldsymbol{f}(\boldsymbol{x,TM})$, where each $\boldsymbol{f}(\boldsymbol{x,Tj})$ denotes the performance metric on task $\boldsymbol{Tj}$ for configuration $\boldsymbol{x}$.
2. Objective Evaluation and Feedback: For each merged model produced by a given configuration, task-specific performance metrics are calculated to assess its effectiveness across tasks. These metrics serve as feedback for refining configurations. This feedback loop enables the optimization algorithm to guide the search towards configurations that better capture the strengths of individual models, enhancing overall performance.
3. Iterative Optimization Process: The multi-objective optimization approach iteratively generates new merging configurations based on feedback from previous evaluations. This process repeats until a preset iteration limit is reached, ensuring that the final configuration achieves a balanced performance across all objectives. The outcome is a merging configuration that maximizes task performance while maintaining desirable generalization characteristics.
By formalizing model merging as a multi-objective optimization problem, we establish a structured foundation for an automated search for effective configurations, ultimately producing a robust, high-performing merged model.
In addition, formalizing the configuration search for model merging as a multi-objective optimization problem presents the following three challenges:
1. How to generate high-quality configurations? To address this challenge, we developed a weak-to-strong method to refine newly generated configurations (Subsection IV-D1), thereby increasing the probability of generating high-quality model merging configurations.
2. How to identify superior configurations with fewer evaluation iterations? To address this challenge, we proposed a filtering approach based on Fisher information to select solutions for evaluation from each generation of candidates (Subsection IV-D2). This method increases the possibility of discovering superior model merging configurations within the constraints of limited evaluation iterations.
3. How to prevent the final merged model from overfitting to specific types of validation tasks? To address this challenge, we introduced an additional sparsity optimization objective, which effectively helps to prevent overfitting on particular evaluation tasks (Subsection IV-E). This objective ensures that the merged model maintains robust overall performance and improved generalization capabilities.
IV Methodology
IV-A Overview
@@
Here, we provide an overview of our proposed automated model merging approach, MM-MO. Our approach is designed to merge multiple pre-trained models into a single, unified model that retains the strengths of each individual model and surpasses the performance of any single model. In Fig. 3, we show the overall process of model merging based on multi-objective optimization, which is divided into three parts:
On the left is a schematic diagram of the automatic search process for model merging configurations. Our method starts with a set of source models, each obtained by performing SFT on different datasets using the same pre-trained backbone. These models have their own advantages for different task types. We merge these models to obtain a merged model, evaluate the performance of the merged model on different tasks as optimization objectives , and calculate the sparsity metric of the merged model as an additional optimization objective. For simplicity, we will denote $\boldsymbol{f}(\boldsymbol{x,Tj})$ later as $\boldsymbol{f_{j}}(\boldsymbol{x})$, $j\in\{1,...,M-1\}$. MM-MO generates new model merging configurations $\boldsymbol{X}=\{\boldsymbol{x}_{1},\boldsymbol{x}_{2},...\}$ based on these objective values, allowing the creation of new merged models. This process is repeated until a preset evaluation iteration limit is reached, ultimately resulting in an optimal model merging configuration.
In the middle is the complete process of using multi-objective Bayesian optimization to search for model merging configurations, detailed in Subsection IV-B and IV-C.
On the right is the step of generating new configurations. We enhanced the acquisition strategy of multi-objective Bayesian optimization by designing a weak-to-strong method to adjust candidate configurations, and we used Fisher information to select the next generation of model merging configurations, detailed in Subsection IV-D.
IV-B Model Merging via Multi-objective Optimization
@@
Our approach builds upon TIES-Merging with DARE [27, 17], utilizing qEHVI to search for the optimal model merging configuration. Since TIES-Merging with DARE first applies sparsification to each source model before merging, our model merging configuration specifically includes the sparsity level of each source model (Density) and the weight proportion of each source model during merging (Weight). The integration of qEHVI allows for an automated and systematic exploration of the parameter space, identifying the best configurations that balance various performance metrics. Here are the detailed steps of our methodology.
1. Initialization: We start with a collection of pre-trained models, each fine-tuned on different tasks. These models and their configurations serve as the initial candidates for merging.
2. Objective Definition: We define multiple objectives that reflect the performance metrics we aim to optimize. Specifically, these objectives include task-specific accuracies, such as those from C-EVAL [31] and GSM8K [32], as well as a sparsity metric $\boldsymbol{g}(\boldsymbol{x})$ used in our experiments. These metrics are crucial as they ensure the merged model performs well across different types of tasks.
3. Surrogate Model Training: qEHVI leverages surrogate models to approximate the objective functions. We train these surrogate models on the initial set of model configurations and their corresponding performance metrics. This step involves creating a probabilistic model that predicts the performance of unseen configurations.
4. Acquisition Function Optimization: The key component of qEHVI is the Expected Hypervolume Improvement (EHVI) acquisition function. This function guides the search by quantifying the expected improvement in the objective space. We optimize this acquisition function to select the most promising model configurations to evaluate next. Specifically, the hypervolume (HV) [33] is a measure of the space dominated by a set of solutions in the objective space, bounded from below by a reference point. The HV indicator is defined as follows:
@@
where $\tilde{f}_{t}\sim p(\boldsymbol{f}|\mathcal{D})$ for $t=1,...,N$ and $\mathcal{X}_{\mathrm{cand}}=\{x_{i}\}_{i=1}^{q}$ [30]. After obtaining candidate configurations using the acquisition function, we adjusted the candidate configurations using the weak-to-strong method. We then selected the configurations based on their Fisher information to ultimately obtain a new generation of model merging configurations. For details, please refer to Subsection IV-D.
5. Parallel Evaluation: Unlike traditional methods, qEHVI supports parallel evaluations, which significantly speeds up the optimization process. In each iteration, we select a batch of model configurations and evaluate their performance on the defined objectives (e.g., C-EVAL and GSM8K accuracy).
6. Updating the Surrogate Model: After evaluating the selected configurations, we update the surrogate models with the new data. This iterative process continues, refining the surrogate models and improving the acquisition function’s accuracy in identifying high-potential configurations.
7. Convergence and Final Selection: The optimization process continues until convergence criteria are met, such as a maximum number of iterations or negligible improvement in the objective space (in subsequent experiments, unless otherwise specified, the termination criterion is set as 30 real evaluations, i.e., FE=30). The final model merging configuration is selected based on the best trade-offs between the multiple performance metrics.
@@
IV-C Pseudocode of MM-MO
Output : final configuration $\boldsymbol{X}_{Final}$
1 Initialize $N$ configurations $\boldsymbol{X}_{0}$ randomly
3 for $k=0$ to $K-1$ do
8 end for
Algorithm 1 The framework of MM-MO
Input : evaluated configurations $\{\boldsymbol{X},\boldsymbol{Y}\}$, acquisition function $\boldsymbol{qEHVI}$, surrogate model $GP$
Output : next-generation configuration $\boldsymbol{X}_{New}$
1 $\boldsymbol{X}_{C}\leftarrow$ Generate candidate configurations based on $\boldsymbol{qEHVI}$ and $\{\boldsymbol{X},\boldsymbol{Y}\}$
2 $\boldsymbol{X}_{H}\leftarrow$ Select high-quality configurations based on $\{\boldsymbol{X},\boldsymbol{Y}\}$
Algorithm 2 Generate new configuration
  1. 1.
    Initialization Phase: The process begins by randomly generating an initial set of $N$ configurations, denoted as $\boldsymbol{X}_{0}$, where $N$ defaults to 10. These configurations are evaluated using the evaluation function $\boldsymbol{f}(\boldsymbol{x})$ and the sparsity metric calculating function $\boldsymbol{g}(\boldsymbol{x})$. The results are combined into an initial set of objective values, $\boldsymbol{Y}_{0}$, which includes both evaluation results and sparsity metrics.
  2. 2.
    Iterative Optimization Phase: The algorithm iterates $K$ times, where $K$ defaults to 4. In each iteration, the current set of configurations $\boldsymbol{X}_{k}$ and their associated objective values $\boldsymbol{Y}_{k}$ are used to train surrogate models (Gaussian Processes, GP) for each objective, which is a commonly used surrogate model in Bayesian optimization. Based on these surrogate models and the current evaluated configurations, a batch of new candidate configurations ${\boldsymbol{X}}_{k}^{B}$ is generated. We set the batch size to 5 in our experiment. The newly generated configurations are added to the existing set, and their objective and sparsity metrics are evaluated. This updates both the configuration set and their objective values for the next iteration.
  3. 3.
    Final Selection Phase: After completing all iterations, the configuration with the highest objective values is selected as the final configuration $\boldsymbol{X}_{Final}$, which serves as the output of the algorithm.
Additionally, when generating new configurations, the algorithm first generates candidate configurations based on the acquisition function (e.g., qEHVI [30]) and the previously evaluated configurations. It then selects high-quality configurations based on objective values, applies a weak-to-strong adjustment method, and finally selects half of the configurations according to the Fisher information for the next generation. This approach can increase the chances of discovering excellent model merging configurations within limited number of evaluations.
IV-D Enhanced Acquisition Strategy
To achieve high-quality model merging configurations within limited number of evaluation iterations, we have further improved the acquisition function of multi-objective Bayesian optimization for the model merging scenario.